OpenStudy (leozap1):

Anyone know of a site that could answer these types of questions? Mathway won't answer these types. *Fan & Medal* Let f(x) = -5x - 4 and g(x) = 6x - 7. Find f(x) + g(x)

OpenStudy (leozap1):

That's a different question, but I'll ask on the site. Thanks.

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{f(x) = -5x - 4} \quad\text{and}\quad \color{orangered}{g(x) = 6x - 7}\]Then,\[\large\rm \color{royalblue}{f(x)}+\color{orangered}{g(x)}=\color{royalblue}{-5x-4}+\color{orangered}{6x-7}\]

zepdrix (zepdrix):

Combine like-terms: -5x + 6x = ? -4 -7 = ?

OpenStudy (leozap1):

-5x + 6x = 1x so just x. -4 - 7 = -11. So the answer is x - 11?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (leozap1):

Yay, thank you! :D

OpenStudy (leozap1):

Could you help me with some more? f(x) = 3x+2 and g(x) = 6x - 7. I tried working it out and I got 9x-5, but that's not an answer choice.

OpenStudy (leozap1):

Nvm It's -3x + 9

OpenStudy (leozap1):

It was negative not positive.